OpenStudy (anonymous):

The equation shows the relationship between x and y: y = −2x + 11 What is the slope of the equation? −11 −2 9 11

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

@Michele_Laino

alones (alones):

Let's solve for x. y=−2x+11 Step 1: Flip the equation. −2x+11=y Step 2: Add -11 to both sides. −2x+11+−11=y+−11 −2x=y−11 Step 3: Divide both sides by -2. −2x−2=y−11−2 x=−12y+112 Answer: x=−12y+112

OpenStudy (anonymous):

whoa cool thanks could you help with some more? :D

alones (alones):

sureee

OpenStudy (anonymous):

kk